OpenStudy (anonymous):

using substituition method x-y=5 3x+4y=13

hero (hero):

3x + 4y = 13 y = x - 5 Substitute x - 5 in place of y in the first equation: 3x + 4(x - 5) = 13 Solve for x

OpenStudy (anonymous):

thank you